For wisdom and a sense of justice to all the world’s leaders, especially now as representatives from the Middle East are attempting to settle their differences with one another…

For a love and respect of this planet, that all may use its abundance wisely and in accordance with your wishes.

For health and healing to those who are sick or dispairing… For the lonely and the friendless, that they may find acceptance… For this broken world, that it may regain a sense of community and mutual love and acceptance… As the whether turns cold, may all those who are poor, needy, or homeless find comfort and succor… We thank you for all of our creature comforts; may we always remember those who are less fortunate

than us… We thank you for the seasons; may we always see the wonder of your creation…